OpenStudy (anonymous):

2 3/4 divided by (-1 4/5) ((Algebra 1))

OpenStudy (anonymous):

-1.52

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

do you want the answer of \[2\frac{ 3 }{ 4 }\div -1\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first look at this \[a \frac{ b }{ c }=\frac{ ac+b }{ c }\]

OpenStudy (ahsome):

Your problem is: \[2\frac{3}{4}÷-1\frac{4}{5}\] First, make them improper fractions by multiplying the denominator with the outside number \[\frac{11}{4}÷\frac{-1}{5}\] Then, make the commong denominators: \[\frac{11*5}{4*5}÷\frac{-1*4}{5*4}\] \[\frac{55}{20}÷\frac{-4}{20}\] Solve from there

OpenStudy (campbell_st):

change both mixed numbers to improper fractions \[\frac{11}{4} \div \frac{-9}{5}\] the rule for dividing by fraction is flip the 2nd fraction and multiply \[\frac{11}{4} \times \frac{5}{-9} \] hope it helps

OpenStudy (ahsome):

^^ That is WAY Easier ^^